Work process of Substitution x=4y+11
4y-5x=9
and x+y=3
x= 6-4y I don't get the work process. and I don't have time for tutoring.. I need help.

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

What sort of help do you need?

Answer by stanbon(75887)

Work process of Substitution
x=4y+11
4y-5x=9
-------------
substitute for "x" and solve for "y":
4y - 5(4y+11) = 9
5y - 20y - 55 = 9
-15y = 64
y = -64/15
---
Substitute into x = 4y+11 to solve for "x":
x = 4(-64/15)+11 = -256/15 + 165/15 = -99/15 = -33/5
===========================================================
x+y=3
x= 6-4y
Substitute for "x" and solve for "y":
6-4y + y = 3
-3y = -3
y = 1
----
Substitute into x = 6-4y to solve for "x":
x = 6-4*1 = 2
===================
Cheers,
Stan H.